# ESG Rating Disagreement and Corporate Total Factor Productivity:Inference and Prediction

### v4

@@ 1 Introduction
Since 1987, the United Nations has reinforced its commitment to sustainable development, and introduced the concept of ESG (Environmental, Social, and Governance) by then-UN Secretary-General Kofi Annan in 2004. Since then, ESG principles have been increasingly integrated into global investment strategies and corporate assessments, with their specific meanings evolving over time. In recent years, China is prominently advocating the philosophy that "green mountains and clear waters are as valuable as mountains of gold and silver", and promoting sustainable development both domestically and internationally. On December 27, 2023, the Guidelines for the Comprehensive Advancement of Beautiful China Construction was issued by the Central Committee of the Communist Party of China and the State Council clearly clarified the exploration and development of ESG assessment frameworks, which shows a strong will of establishing ESG Evaluation System with Chinese Characteristics.
Despite the implementation of various ESG policies in recent year, there are still significant disagreement in ESG rating system. From the moment the ESG concept was introduced in China, multiple rating agents have independently published their ESG rating reports. However, these reports often display significant discrepancies in ratings, which undoubtedly has a substantial impact on the companies being evaluated. Examining the economic consequences of these ESG rating disagreements is of considerable value for both enterprises and regulatory bodies. The main recent studies are: Wang et al. (2024); Zhao et al. (2024) argue that disagreements in ESG ratings are significantly negatively correlated with stock returns; ESG rating disagreement increase analyst forecast errors(Liu et al., 2024); ESG rating disagreement drives firms towards green innovation (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); Moreover, increased disagreement in ESG ratings leads to higher levels of real earnings management both in the current period and over the next one to two years (Li and Cheng, 2024); The firms with agreed high ESG rating can benefit through cost reduction and larger amount debt-financing(Guo et al., 2024); Besides,the disagreement in ESG ratings may poses a potential risk of stock price collapse(Luo et al., 2023); What’s more, ESG rating disagreement significantly increases audit fees due to information asymmetry(Ling et al., 2024).
However, the focus of most recent studies has been less on exploring the impact of ESG rating disagreements on the overall development of a company. Therefore, we select TFP as the subject of our analysis in this study, as it serves as an indicator of a company’s high-quality development.
@@ 2 Theoretical Analysis and Research Hypotheses
According to information asymmetry theory, ESG rating disagreements can confuse stakeholders and contribute to the degree of information asymmetry between companies and investors. ESG ratings, as non-financial information, provide stakeholders with a more comprehensive insight into companies, thereby alleviating information asymmetry(Xiaohong and Zhenghan, 2023). Conversely, significant ESG rating disagreements can intensify information asymmetry, creating a noise effect that leads to investor misjudgments(Berg et al., 2022). Thereby, increasing financing constraints and ultimately resulting in a decline in TFP. Simultaneously, based on stakeholder theory, companies may take certain remedial measures in response to ESG rating disagreements to balance the interests of various stakeholders. Companies might intensify green innovation under the stimulus of ESG rating disagreements(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); However, as this initiative cannot cater to all stakeholders under noise effect caused by ESG rating disagreement, it may lead some groups to perceive "greenwashing" behaviors, which ultimately undermines the company’s financing efforts and results in a decline in TFP(Peng and Xie, 2024). Meanwhile, the divergence in ESG ratings may also drive companies to focus on improving their ESG recognition, potentially causing an imbalance between the interests of external investors and internal employees. This imbalance could lead to a human capital outflow, ultimately affecting the development of TFP.
In summary, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
@@ 3.1 Sample Selection
We utilize financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., alongside ESG ratings from four agencies: Huazheng, Wind, SynTao Green Finance, and Susallwave. The data processing procedures follow these steps: (1) excluding companies in the financial and real estate industries; (2) excluding ST, *ST, and delisted companies; (3) excluding companies with only one ESG rating in a given year; (4) excluding companies with missing key indicators; and (5) applying Winsorization at the 1% level on both tails. All data is sourced from the CSMAR database, the CNRDS database, and the four rating agencies, resulting in 3656 firms and 13417 samples in baseline regression.
@@ 3.3 Variable Descriptions
Table 1: Definition of Main Variables
Symbol Source Description
$TFP$ CSAMR Estimated using the LP, OP and GMM method
$Dis$ ESG Rating Agencies Calculated based on the method described above
$TobinQ$ CSAMR Tobin’s Q ratio
$Top10$ CSAMR Shareholding of top 10 shareholders / total shares
$ListAge$ CSAMR Natural logarithm of years since the firm’s IPO
$Lev$ CSAMR Total liabilities at year-end / total assets at year-end
$Size$ CSAMR Natural logarithm of total assets at year-end
$Avg$ CSAMR Natural logarithm of operating income / number of employees
$Der$ CSAMR Total liabilities at year-end / total equity at year-end
$EI$ CNRDS Natural logarithm of the number of green patents obtained in a given year
$FC$ CSAMR WW and KZ indices, the larger they are the more constrained the financing is
$HC$ CSAMR Natural logarithm of number of graduate students and above and natural logarithm of number of undergraduate students and above
@@ 4.1 Baseline Regression
Table 2: Baseline Regression Results
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0675*** -0.0536** -0.0536** -0.0447*** -0.0357** -0.0373**
(-2.7452) (-2.1838) (-2.2098) (-2.9449) (-2.2503) (-2.2395)
$TobinQ$0.0032 -0.0017 -0.0007
(-0.5471) (-0.3953) (-0.1795)
$Top10$0.0010 -0.0009 -0.0005
(-0.9976) (-0.9496) (-0.4838)
$ListAge$ 0.0989** 0.0855** 0.0160
(2.4952) (2.3667) (0.4241)
$Lev$0.1697 -0.1557 -0.1481
(-1.5121) (-1.5361) (-1.4348)
$Size$ 0.3737*** 0.2581*** 0.2245***
(13.129) (14.214) (10.939)
$Avg$ 0.2292*** 0.2695*** 0.2671***
(2.7852) (2.8274) (2.7426)
$Der$$0*$0190* -0.0211**
(-2.0405) (-1.8342) (-1.9816)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
Doubel Clustering YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0015 0.0011 0.0011 0.1785 0.1670 0.1409
  • Note: t-statistics are in parentheses. * p<0.1, ** p<0.05, *** p<0.01. The same applies below.
@@ 4.2.1 The Role of Green Innovation
Table 3 and figure1 reports the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. The result seem contradict exiting researches, which argue ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
Table 3: Green Innovation Mechanism
Variables TFP
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$0.0264 -0.0205 -0.0199 -0.0167 -0.0130 -0.0136
(-1.2045) (-1.0114) (-0.9590) (-0.8367) (-0.7014) (-0.7069)
$EI$ 0.0412*** 0.0283*** 0.0253*** 0.0223*** 0.0142*** 0.0126***
(8.4064) (6.2383) (5.4522) (4.9596) (3.3848) (2.8967)
$Dis*EI$0.0366*** -0.0292*** -0.0296*** -0.0249** -0.0199** -0.0207**
(-3.1706) (-2.7348) (-2.7184) (-2.3693) (-2.0401) (-2.0432)
$TobinQ$0.0030 -0.0016 -0.0007
(-0.7895) (-0.4611) (-0.1778)
$Top10$$0$$0$0005
(-1.7413) (-1.7832) (-0.8682)
$ListAge$ 0.1012*** 0.0869*** 0.0172
(3.4856) (3.2139) (0.6122)
$Lev$0.1692*** -0.1550*** -0.1472***
(-3.2834) (-3.2318) (-2.9594)
$Size$ 0.3664*** 0.2537*** 0.2208***
(28.243) (21.017) (17.632)
$Avg$ 0.2297*** 0.2698*** 0.2673***
(22.666) (28.608) (27.331)
$Der$$0**$0190*** -0.0211***
(-3.1573) (-3.0646) (-3.2805)
Controls NO NO NO YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
$N$ 13417 13417 13417 13417 13417 13417
$R^{2}$ 0.0092 0.0052 0.0042 0.1806 0.1680 0.1417
Figure 1: Marginal Effect
To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the tabel 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
Table 4: Heterogeneity of Green Innovation Mechanism
Variables TFP
High $Dis$ Low $Dis$
$TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$ $TFP_{LP}$ $TFP_{OP}$ $TFP_{GMM}$
$Dis$ 0.0527 0.0616 0.0501 -0.2268*** -0.1532** -0.1554**
(1.2360) (1.5631) (1.2368) (-2.8516) (-2.0471) (-1.9903)
$EI$ 0.0364*** 0.0280*** 0.0230** 0.0092 0.0043 0.0041
(3.1495) (2.6193) (2.0914) (1.4601) (0.7258) (0.6628)
$Dis*EI$0.0522** -0.0461** -0.0381** 0.0433 0.0280 0.0260
(-2.1556) (-2.0558) (-1.6525) (1.2832) (0.8798) (0.7855)
$TobinQ$0.0009 -0.0012 0.0020 -0.0025 -0.0020 -0.0013
(-0.1395) (0.1681) (0.3447) (-0.4475) (-0.3853) (-0.2318)
$Top10$0.0008 -0.0012 -0.0005 -0.0001 0.0004 0.0008
(-0.8389) (-1.3029) (-0.5405) (-0.1141) (0.4905) (0.8934)
$ListAge$ 0.1137** 0.0833* 0.0257 0.0609 0.0529 -0.0332
(2.2567) (1.7886) (0.5359) (1.4289) (1.3186) (-0.7922)
$Lev$0.1140 -0.0997 -0.0976 -0.2786*** -0.2703*** -0.2712***
(0.1716) (-1.2930) (-1.2306) (-3.2945) (-3.3962) (-3.2670)
$Size$ 0.3607*** 0.2525*** 0.2173*** 0.3523*** 0.2355*** 0.2123***
(16.816) (12.732) (10.655) (17.502) (12.434) (10.746)
$Avg$ 0.2126** 0.2424*** 0.2447*** 0.2301*** 0.2706*** 0.2651***
(13.198) (16.277) (15.983) (13.969) (17.451) (16.389)
$Der$ 0.0224** -$0*$0226** 0.0046 0.0052 0.0036
(-2.2377) (-2.2735) (-2.3751) (0.3638) (0.4404) (0.2910)
Controls YES YES YES YES YES YES
Year YES YES YES YES YES YES
Id YES YES YES YES YES YES
$N$ 6774 6774 6774 6643 6643 6643
$R^{2}$ 0.1578 0.1434 0.1229 0.1717 0.1533 0.1310
4.2.2 The Role of Financing Constraints
@@
Table 5 reports the regression results. The $Disb$ is the variable by one period. So we concludes that increased ESG rating disagreement reduces corporate financing constraints, thereby lowering TFP.
Table 5: Financing Constraints Mechanism
Variables Financing Constraints
$WW$ $WW$ $KZ$ $KZ$ $WW$ $WW$ $KZ$ $KZ$
$Dis$ 0.0122*** 0.2166*** 0.0089*** 0.1815***
(4.7408) (3.0974) (3.6730) (2.6598)
$Disb$ 0.0122*** 0.2142*** 0.0074*** 0.1392**
(4.7691) (3.0604) (3.3430) (2.2547)
$TobinQ$0.0022*** -0.0006 0.2724*** 0.0237
(-3.9269) (-0.9530) (19.528) (1.4433)
$Top10$0.0007*** -0.0006*** 0.0011 -0.0061**
(-8.8614) (-6.9840) (0.4783) (-2.3496)
$ListAge$ 0.0204*** 0.0143*** 0.5530*** 1.2048***
(6.1906) (2.9544) (6.5333) (9.5882)
$Lev$ 0.0370*** 0.0359*** 6.7446*** 3.1453***
(5.0820) (4.5416) (32.842) (14.109)
$Size$0.0614*** -0.0455*** -0.4512*** -0.2274***
(-36.832) (-22.973) (-9.5683) (-4.0781)
$Avg$$0**$0113*** -0.5288*** -0.0107
(-18.835) (-7.3282) (-13.531) (-0.2432)
$Der$ 0.0030*** 0.0011 0.0191 -0.0027
(3.4646) (1.2994) (0.6820) (-0.0931)
Controls NO NO NO NO YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
$N$ 10059 10059 13417 13417 10059 10059 13417 13417
$R^{2}$ 0.0034 0.0034 0.0010 0.0010 0.2647 0.1221 0.2250 0.0517
  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and financing constraints in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and financing constraints in year t.
@@ 4.2.3 The Role of Human Capital
Table 6: Human Capital Mechanism
Variables Human Capital
$Master$ $Master$ $Undergrad$ $Undergrad$ $Master$ $Master$ $Undergrad$ $Undergrad$
$Dis$0.0595*** -0.0526*** -0.0345** -0.0316**
(-2.8695) (-3.1495) (-1.7625) (-2.0461)
$Disb$0.0598*** -0.0528** -0.0322* -0.0248*
(-2.8867) (-3.1608) (-1.8268) (-1.9302)
$TobinQ$ 0.0197*** -0.0067 0.2724*** -0.0036
(5.0228) (-1.4155) (19.528) (-0.9603)
$Top10$ 0.0029*** 0.0008 0.0011 0.0003
(4.4620) (1.0079) (0.4783) (0.4613)
$ListAge$0.0010 0.1365*** 0.5530*** 0.1578***
(-0.0413) (3.7156) (6.5333) (5.5267)
$Lev$ 0.0536 -0.0620 6.7446*** 0.0894*
(0.8913) (-0.9460) (32.842) (1.7679)
$Size$ 0.6722*** 0.4376*** -0.4512*** 0.4410***
(49.844) (27.508) (-9.5683) (35.103)
$Avg$0.1597*** -0.0907*** -0.5288*** -0.0739***
(-14.054) (-7.0579) (-13.531) (-7.5015)
$Der$0.0256*** -0.0128 0.0191 -0.0184***
(-3.1008) (-1.5259) (0.6820) (-2.8025)
Controls NO NO NO NO YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
$N$ 10461 10461 12595 12595 10461 10461 12595 12595
$R^{2}$ 0.0011 0.0011 0.0010 0.0011 0.2790 0.1145 0.4112 0.1511
  • Note: Similar to the robustness tests, $Dis$ corresponds to regressions on ESG rating disagreement in year t and human capital in year t+1, while $Disb$ corresponds to regressions on ESG rating disagreement in year t-1 and human capital in year t.
@@ 4.3 Heterogeneity Analysis
To further explore whether the dampening effect of ESG rating disagreement on TFP exhibits heterogeneity, this study conducts a subgroup regression analysis. Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data was Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
@@
Table 7: Heterogeneity Analysis
Variables TFP
State-Owned Non-State-Owned Capital-Intensive Non-Capital-Intensive High-Pollution Low-Pollution High-Tech Non-High-Tech
$TFP_{LP}$
$Dis$0.0622** -$0$0322 -0.0462*** -0.0239 -0.0481*** -0.0425** -0.0043
(-2.0516) (-1.8294) (-1.1009) (-2.8884) (-1.0401) (-2.7959) (-2.2163) (-0.1653)
Chow Test 500.19*** 289.12*** 544.57*** 756.47***
$R^{2}$ 0.1573 0.1886 0.1741 0.1817 0.2010 0.1754 0.1751 0.1570
$TFP_{OP}$
$Dis$0.0553* -0.0268 -0.0097 -0.0458*** -0.0070 -0.0493** -0.0365** -0.0048
(-1.8608) (-1.5208) (-0.3229) (-2.6338) (-0.3149) (-2.4735) (-2.2763) (-0.1639)
Chow Test 529.58*** 269.31*** 534.66*** 760.50***
$R^{2}$ 0.1759 0.1593 0.1672 0.1663 0.1930 0.1595 0.1561 0.1526
$TFP_{GMM}$
$Dis$0.0618** -$0$0278 -0.0471*** -0.0134 -0.0519** -0.0413** -0.0122
(-2.1381) (-1.6745) (-0.9906) (-2.6889) (-0.6185) (-2.5484) (-2.4165) (-0.4205)
Chow Test 498.97*** 248.15*** 475.42*** 698.54***
$R^{2}$ 0.1607 0.1254 0.1504 0.1377 0.1691 0.1321 0.1256 0.1315
Controls YES YES YES YES YES YES YES YES
Year YES YES YES YES YES YES YES YES
Id YES YES YES YES YES YES YES YES
Double Clustering YES YES YES YES YES YES YES YES
$N$ 4416 9001 2342 11075 3900 9517 8454 4963
@@ 5 Further Discussion
In the above analysis we can get that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression. In order to further explore the relationship between ESG rating divergence and TFP under the nonlinear assumption, we consider a regression analysis using a machine learning approach that can be interpreted with SHapley Additive exPlanations(SHAP) introduced by Scott et al. (2017), which can observe whether ESG rating disagreement is crucial for explaining TFP under the assumption of nonlinear system and explore the nonlinear influence relationship between the two variables.
Following Xue et al. (2024), we use the XGBoost machine learning model introduced by Chen and Guestrin (2016) to predict $TFP_{LP}$. XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in some tasks. In this paper, the Optuna library developed by Akiba et al. (2019), combed with early stopping and cross-validation strategies is employed for parameter selection. After that, we adopt the SHAP (SHapley Additive exPlanations) method proposed by Scott et al. (2017) to explain XGBoost. The variables include the ESG rating disagreement ($Dis$ $Disb$), all control variables, and institutional ownership percentage ($Inst$).
@@
Table 8: Performance Metrics for Training and Test Sets
Metric Training Set 1 Training Set 2 Training Set 3 Training Set 4 Test Set 1 Test Set 2 Test Set 3 Test Set 4
MSE 0.2609 0.2274 0.2818 0.4714 0.3492 0.3300 0.3835 0.4728
RMSE 0.5108 0.4769 0.5309 0.6866 0.5909 0.5744 0.6192 0.6876
MAE 0.3995 0.3729 0.4145 0.5383 0.4589 0.4523 0.4818 0.5415
MAPE 4.7786% 4.4569% 4.9457% 6.4047% 5.4649% 5.3644% 5.7293% 6.4399%
R2 0.7685 0.7969 0.7500 0.5794 0.7045 0.7176 0.6755 0.5955
  • Note: Set1 is the XGBoost performance of $Dis$ in the same period as the TFP, Set2 is the XGBoost performance of $Dis$ in the preceding period, Set3 is the XGBoost performance excluding $Dis$, and Set4 is the model performance of the linear model (no fixed effects) and $Dis$ in the preceding period.
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we applied SHAP to interpret the results of the XGBoost model. SHAP values are grounded in cooperative game theory and provide a way to understand each feature’s contribution to the model output by treating features as "players" in a game. The contribution, or Shapley value $\phi_{i}$, for a particular feature $i$ is computed by averaging its marginal contributions across all possible combinations of features.
@@
In this equation:
  • $N$ represents the set of all features, while $S$ is a subset of $N$ that does not include feature $i$.
  • $|S|$ denotes the size of subset $S$.
  • $f(S)$ is the model output based on the subset of features $S$.
  • $f(S\cup\{i\})$ is the output when feature $i$ is added to $S$.
@@
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP [^3] The importance of each feature is calculated as the percentage of the average of the absolute SHAP values from each to the mean of the absolute SHAP values of all features; in short, a higher percentage means that the feature is more important among all features.. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task [^4] The beeswarm plot shows the effect of different values of feature on TFP, the color of the dots in the plot shifts toward red for larger eigenvalues and toward blue for smaller eigenvalues and points positioned on the horizontal axis with values less than 0 indicate a negative effect on the dependent variable in the model, while points with values greater than 0 signify a positive effect.. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
Figure 2: Feature Importance
Figure 3: SHAP Beeswarm Plot

### v5

@@ 1 Introduction
Since 1987, the United Nations has championed sustainable development and introduced ESG (Environmental, Social, and Governance) in 2004. Over time, these principles have become key components of global investment strategies and corporate assessments. In recent years, China has embraced sustainable development through initiatives like the philosophy “green mountains and clear waters are as valuable as mountains of gold and silver.” On December 27, 2023, the Central Committee of the Communist Party of China and the State Council issued the Guidelines for the Comprehensive Advancement of Beautiful China Construction, which emphasized creating an ESG Evaluation System with Chinese Characteristics.
Despite recent ESG policy efforts, significant disagreements persist within ESG rating systems. Since ESG ratings were introduced in China, multiple agencies have issued their own reports, often resulting in substantial discrepancies. Such differences have considerable implications for evaluated companies. Recent studies highlight the economic impacts of ESG rating disagreement: stock returns suffer (Wang et al., 2024; Zhao et al., 2024); analyst forecast errors rise (Liu et al., 2024); green innovation encouragement (Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a); real earnings management increases in both the short and long term (Li and Cheng, 2024); firms with consistently high ratings benefit from lower costs and increased debt financing (Guo et al., 2024); disagreement poses a risk of stock price collapse (Luo et al., 2023); audit fees increase due to information asymmetry (Ling et al., 2024); and market pressures on managers intensify (Chen and Wang, 2025).
However, most recent studies have not sufficiently examined ESG rating disagreement’s impact on a company’s overall development. To address this gap, we focus on TFP as an indicator of high-quality development in this study.
@@ 2 Theoretical Analysis and Research Hypotheses
ESG rating disagreement can confuse stakeholders and increase information asymmetry between companies and investors. While ESG ratings help reduce information asymmetry by providing non-financial insights(Xiaohong and Zhenghan, 2023), disagreement can amplify asymmetry, creating noise that misleads investors(Berg et al., 2022), raises financing constraints, and ultimately lowers TFP. Stakeholder theory suggests that companies may respond to disagreement by pursuing green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024a), but this can appear as “greenwashing” to some stakeholders, undermining financing and reducing TFP(Peng and Xie, 2024). Disagreement may also prompt firms to prioritize ESG recognition, causing tension between external investors and internal employees, potentially leading to a human capital outflow and further TFP decline.
Hence, we propose the research hypothesis H1: ESG rating disagreement reduces corporate TFP.
@@ 3.1 Sample Selection
We use financial data of A-share listed companies in China from 2015 to 2022111Prior 2015, the widely recognized local ESG rating agency in China was Huazheng., along with ESG ratings from Huazheng, Wind, SynTao Green Finance, and Susallwave. The data processing steps are as follows: (1) exclude companies in the financial and real estate sectors; (2) exclude ST, *ST, and delisted firms; (3) exclude companies with only one ESG rating in a year; (4) exclude firms with missing key indicators; (5) apply Winsorization at the 1% level on both tails. The data comes from the CSMAR and CNRDS databases and the four rating agencies, yielding 3656 firms and 13,417 samples in the baseline regression.
@@ 3.3 Variable Descriptions
[Insert Table 1 Here]
@@ 4.1 Baseline Regression
[Insert Table 2 Here]
@@ 4.2.1 The Role of Green Innovation
Table 3 and Figure 1 report the results. The results show that when there are no ESG rating discrepancies, green innovation promotes the improvement of TFP. Conversely, when there are ESG rating discrepancies, it strengthens the dampening effect of ESG rating disagreement on TFP. The result seem contradict exiting researches, which argue ESG rating disagreement can drive firms to enhance green innovation(Hou and Xie, 2024; Geng et al., 2024; Chen et al., 2024b). This observation is very likely caused by a remedial action taken by firms in response to ESG rating disagreement, as Hou and Xie (2024) mentioned, such ESG rating disagreement is negatively correlated with the quality of corporate green innovation. According to stakeholder theory, firm’s remedial green innovations may be seen as "greenwashing" when there is disagreement over ESG ratings, which will lead to higher costs of debt financing for corporations, ultimately bringing about a decline in TFP(Peng and Xie, 2024).
[Insert Table 3 Here]
[Insert Figure 1 Here]
To further investigate whether the moderating effect varies across levels of ESG rating disagreement, we conducted regressions separately for the upper and lower 50% quartiles of ESG rating disagreement, and the Table 4 show that the moderating effect is significant only when the ESG rating disagreement is high, further supporting the notion that substantial ESG rating disagreement may lead to stakeholder distrust in firms’ green innovation.
[Insert Table 4 Here]
4.2.2 The Role of Financing Constraint
@@
Table 5 reports the regression results. The $Disb$ is the variable by one period. So we concludes that increased ESG rating disagreement strengthens corporate financing constraints, thereby lowering TFP.
[Insert Table 5 Here]
@@ 4.2.3 The Role of Human Capital
[Insert Table 6 Here]
@@ 4.3 Heterogeneity Analysis
To further explore whether the dampening effect of ESG rating disagreement on TFP exhibits heterogeneity, this study conducts a subgroup regression analysis. Given that the sample size decreases after subgrouping, making the results more susceptible to the influence of outliers, the data is Winsorized at the 2% level on both ends before performing the subgroup regression and conducting a Chow Test on $Dis$.
@@
[Insert Table 7 Here]
@@ 5 Further Discussion
In the above analysis we can see that ESG rating disagreement brings about a decrease in TFP, which is obtained based on linear regression. In order to further explore the relationship between ESG rating divergence and TFP under the nonlinear assumption, we consider a regression analysis using a machine learning approach that can be interpreted with SHapley Additive exPlanations(SHAP) introduced by Scott et al. (2017), which can observe whether ESG rating disagreement is crucial for explaining TFP under the assumption of nonlinear system and explore the nonlinear influence relationship between the two variables.
Following Xue et al. (2024), we use the XGBoost machine learning model introduced by Chen and Guestrin (2016) to predict $TFP_{LP}$. XGBoost is an ensemble learner based on gradient boosted trees (GBDT) that has demonstrated excellent performance in many applications, even outperforming deep neural networks in tabular data(Grinsztajn et al., 2022). In this paper, the Optuna library developed by Akiba et al. (2019), combed with early stopping and cross-validation strategies is employed for parameter selection. After that, we adopt the SHAP (SHapley Additive exPlanations) method proposed by Scott et al. (2017) to explain XGBoost. The variables include the ESG rating disagreement ($Dis$ $Disb$), all control variables, and institutional ownership percentage ($Inst$).
@@
[Insert Table 8 Here]
To analyze the influence of ESG rating disagreement on $TFP_{LP}$ in our predictive task, we use SHAP to interpret the results of the XGBoost model. SHAP values are grounded in cooperative game theory and provide a way to understand each feature’s contribution to the model output by treating features as "players" in a game. The contribution, or Shapley value $\phi_{i}$, for a particular feature $i$ is computed by averaging its marginal contributions across all possible combinations of features.
@@
where $N$ represents the set of all features, $S\subseteq N\setminus\{i\}$ is a subset excluding feature $i$, $|S|$ denotes the cardinality of subset $S$, $f(S)$ indicates the model output using feature subset $S$, and $f(S\cup\{i\})$ represents the output when feature $i$ is added to $S$.
@@
In this study, based on the previous training results, we chose the model with one-period lagged ESG rating disagreement for our SHAP analysis, and the SHAP values were calculated for the test set. Figure 2 reports the feature importance, showing that the contribution ratio of $Disb$ is 0.14, again indicating that $Disb$ is crucial for predicting TFP. Figure 3 reports the SHAP beeswarm plot, it can be observed that high ESG rating divergence suppresses a company’s TFP, while in companies with low ESG ratings and small divergences, ESG rating disagreement is even positively correlated with TFP in the prediction task. The reason for this phenomenon is that the investment market is more tolerant of small ESG rating divergence, and listed companies will not face significant financing pressure due to small divergence.
[Insert Figure 2 Here]
[Insert Figure 3 Here]